# Contour Integration Underlies Human-Like Vision

### v1

@@ 2 Related work
Contour integration in DNNs. Classical contour integration was studied in DNNs by Linsley et al. (2018), who showed that traditional convolutional neural networks struggle to learn to integrate line segments across the visual field in an abstract task. Inspired by horizontal connectivity in the primate brain, they developed a horizontal gated recurrent unit network, which was able to learn and solve the contour integration task. (Funke et al., 2021) took a different approach: they fine-tuned an ImageNet-trained ResNet-50 (He et al., 2016) on line segment images with open and closed contours. They found that while the model did well on the task in-distribution, generalization proved catastrophically difficult.
@@ 4 Results
All models were trained for 100 epochs using two A100 GPUs per model, with a total batch size of 512. For more training and data preparation details, see Appendix A.6.
We found that while it is expectedly possible to improve performance on fragmented objects by directly training models on similar types of stimuli, it comes at a cost – unlike previous studies (Geirhos et al., 2018), we were not able to simultaneously improve performance on full color images while also training models to exhibit improved robustness to contour fragmentation (Figure 7). Rather, models that improved performance on fragmented stimuli (IN + Segments, IN + Phosphenes, IN + Segments + Phosphenes) also saw minor drops in full color image recognition (RGB).
@@
5 Discussion
Summary. This work presented a large-scale comparison of human and DNN performance on a contour integration task. By carefully designing an experiment meant to specifically test the failure points of DNN models, we found that across an evaluation of over 1,000 DNNs, they fail to capture human behavioral performance, even when allowing for additional decoder-fitting trials. We found that the root cause of the model failure is an inability to integrate contours as measured by the difference in performance on images with directional segments and directionless phosphenes: models that had a larger integration bias also performed better on our contour integration task in general. Our definition of integration bias builds on a vast body of evidence both in human behavior and neural data that shows the specificity of contour integration to directionally aligned line segments (Bosking et al., 2000, 1997; Roelfsema, 2006; Kovács & Julesz, 1993; Kapadia et al., 1995). We found that integration bias arises from model training on large datasets – models trained on more data exhibited a larger integration bias. Finally, training models on fragmented images led not only to integration bias, but also to a larger shape bias than shape-trained models (Geirhos et al., 2018).
Limitations. In this work, we focused on observational findings. We have made several efforts to remedy this: first, our model set covered much of the extant landscape of open access pre-trained vision models and, combined with models we additionally trained ourselves, constituted what is the largest human-model comparison to date by approximately an order of magnitude. Second, we synthesized several new perturbed versions of ImageNet and trained several models on different combinations of ImageNet and its perturbed counterparts.
Acknowledgements
We thank Yingtian Tang for helpful discussions, and Marc Repnow for technical support.
Impact Statement
Appendix A Appendix
A.1 Integration bias predicts classification accuracy
While we did not evaluate all 1,038 models on ImageNet, we evaluated the same subset on ImageNet as we did on ImageNet-C (Figure 4). We find what we found for accuracy on our task – performance is predicted well by integration bias.
A.2 Human experiment category selection
A.3 Human control experiment
A.4 Further model details
@@
A.5 Additional details on training datasets.
@@
A.6 Training models on fragmented ImageNet
@@
A.7 Model set for joint behavioral analysis
@@
A.8 GPT-4o experimental methodology

### v2

@@ 2 Related work
Contour integration in DNNs. While some of the earliest non-DNN models of contour integration were studied earlier (Li, 1998), classical contour integration was studied in DNNs by Linsley et al. (2018), who showed that traditional convolutional neural networks struggle to learn to integrate line segments across the visual field in an abstract task. Inspired by horizontal connectivity in the primate brain, they developed a horizontal gated recurrent unit network, which was able to learn and solve the contour integration task. (Funke et al., 2021) took a different approach: they fine-tuned an ImageNet-trained ResNet-50 (He et al., 2016) on line segment images with open and closed contours. They found that while the model did well on the task in-distribution, generalization proved catastrophically difficult.
@@ 4 Results
In addition, we tested a Shape-biased model (Geirhos et al., 2018). All of our fragmented object models were trained for 100 epochs using two A100 GPUs per model, with a total batch size of 512. For more training and data preparation details, see Appendix A.9, and for additional controls on contour training, see Appendix Figure 12.
We found that while it is possible to improve performance on fragmented objects by directly training models on similar types of stimuli, it comes at a cost – unlike previous studies (Geirhos et al., 2018), we were not able to simultaneously improve performance on full color images while also training models to exhibit improved robustness to contour fragmentation (Figure 7). Rather, models that improved performance on fragmented stimuli (IN + Segments, IN + Phosphenes, IN + Segments + Phosphenes) also saw minor drops in full color image recognition (RGB). While this is expected, it also paves the way to study contour integration mechanistically in smaller DNN models.
Figure 7: Model performance of models trained on ImageNet-1k and fragmented ImageNets. All models were evaluated zero-shot. Left: Our trained model performance on the fragmented object task. Middle: Individual models’ integration bias. Right: Model shape bias as measured by the cue conflict dataset (Geirhos et al., 2021). Our models trained on phosphenes, and a combination all reach higher shape bias than the Stylized ImageNet-trained model (Geirhos et al., 2018).
5 Discussion
Summary. This work presented a large-scale comparison of human and DNN performance on a contour integration task. By carefully designing an experiment meant to specifically test the failure points of DNN models, we found that across an evaluation of over 1,000 DNNs, they fail to capture human behavioral performance, even when allowing for additional decoder-fitting trials. We found that the root cause of the model failure is an inability to integrate contours as measured by the difference in performance on images with directional segments and directionless phosphenes: models that had a larger integration bias also performed better on our contour integration task in general. Our definition of integration bias builds on a vast body of evidence both in human behavior and neural data that shows the specificity of contour integration to directionally aligned line segments (Bosking et al., 2000, 1997; Roelfsema, 2006; Kovács & Julesz, 1993; Kapadia et al., 1995). We found that integration bias arises from model training on large datasets – models trained on more data exhibited a larger integration bias. Finally, training models on fragmented images led not only to integration bias, but also to a larger shape bias than shape-trained models (Geirhos et al., 2018). This allows for researchers to study the computations underlying human-like integration bias in DNNs, and to obtain a high model shape bias simultaneously with a high integration bias, by training on fragmented objects.
Limitations. In this work, we focused on observational findings. We have made several efforts to remedy this: first, our model set covered much of the extant landscape of open access pre-trained vision models and, combined with models we additionally trained ourselves, constituted what is the largest human-model comparison to date by approximately an order of magnitude. Second, we synthesized new perturbed versions of ImageNet and trained models on different combinations of ImageNet and its perturbed counterparts.
Acknowledgements
We thank Yingtian Tang for helpful discussions, and Marc Repnow for technical support. BL was supported by a grant from the Swiss National Science Foundation (N.320030_176153; “Basics of visual processing: from elements to figures”). ES was supported by a grant from ERA-NET NEURON (Ref Nr: NEURON-051).
Impact Statement
Appendix A Appendix
A.1 Integration bias predicts classification accuracy
We evaluated model performance on ImageNet and found the same results as we did on ImageNet-C (Figure 4). Integration bias predicts performance across tasks, from fragmented stimuli (Figure 6a) to ImageNet ($r=0.67,p<0.01$).
A.2 Human experiment category selection
A.3 Human control experiment
A.4 Human performance scaling across conditions
We analyzed the difference in how performance across the segments and the phosphenes conditions scales. We found little difference (phosphenes vs segments) in humans – the difference in performance in these two cases is mostly a shift of intercept instead of a change in slope across the number of elements, though while the effect of difference in slopes is small, it is still technically significant ($t=2.87,p<0.01$). See Figure 10 for details. This means that the directionality of the segments only slightly affects performance positively over the directionless prosphenes as element density increases.
Figure 10: Human accuracy split by condition (phosphenes: light green; segments: dark green).
A.5 Integration is not predicted by receptive field size alignment
We evaluated the same subset of models as in Figure 6 on two Brain-Score benchmarks that test for the similarity of the effective receptive field size of a model to a primate counterpart. These are the Grating summation field (GSF) and surround diameter (Marques et al., 2021; Cavanaugh et al., 2002) benchmarks measured in Macaque V1. See Figure In short, we do not find a statistically significant relationship between either measure of receptive field size similarity and fragmented object recognition accuracy (GSF $r=0.2655,p=0.0653$; surround diameter $r=0.276,p=0.055$)
Figure 11: Model performance correlated to alignment in two benchmarks: the Grating summation field (GSF) and surround diameter (Marques et al., 2021; Cavanaugh et al., 2002).
A.6 Contour-trained models do not achieve as high shape bias or integration bias
As in Figure 7, we also trained additional control models to investigate whether contour-based models, and the distribution shift in image statistics caused by that, could be at cause for the results in Figure 7. We trained the following models:
IN only: ImageNet-1k baseline with no additional images.
@@ A.6 Contour-trained models do not achieve as high shape bias or integration bias
We found that the effect of training solely on contours is minimal to non-existent, and thus the shift in image statistics does not account for the observed differences in integration and shape biases (Figure 12).
@@
A.7 Further model details
@@
A.8 Additional details on training datasets.
@@
A.9 Training models on fragmented ImageNet
@@
A.10 Model set for joint behavioral analysis
@@
A.11 Stimulus synthesis
We implemented stimulus synthesis using (Rotermund et al., 2024), which implements the following algorithm to filter image $C(x,y,\varphi)$:
[…]
Where
$c_{0}=\exp\!\Bigl{(}-2\bigl{(}\tfrac{\pi\sigma}{\lambda}\bigr{)}^{2}\Bigr{)}.$
The convolution operation is implemented via a 2D Fast Fourier Transform, and $I(x,y)$ is the grayscale image pixel value at $(x,y)$. Thus, the resulting contour image $C(x,y,\varphi)$ contains the value of the local contour of orientation $\varphi$ at position $(x,y)$. We set the spatial filter size to $\sigma=0.06$ arcdeg and the spatial scale to $\lambda=0.12$ arcdeg.
For the placement of phosphenes and segments, we place elements on the contours of the object preferentially depending on the strength of the contour and its directionality. We use this algorithm both for our experimental stimuli, as well as ImageNet-1k images. For ImageNet-1k images, we also perform a background removal using rembg (Gatis, 2023) before applying this algorithm.
A.12 GPT-4o experimental methodology
@@
A.13 Image background color control experiment.
To test whether the specific image color statistics affect model performance or cause difficulties, we tested a small subset of models on red background images. We performed a t-test to test for the difference in the group performance of the fragmented conditions and found that there was no difference in model accuracy under these two conditions ($t=0.323,p=0.749$). This shows that the specific model training distribution is not the reason for model performance here, and that the results are robust against a change of background (Figure 15).
Figure 14: Red background control. (Left): Examples of images with a red, instead of black, background. (Right): Results show no sensitivity to the background color.
A.14 Differences between ViT and Convolutional architectures.
We conducted several analyses comparing transformers and CNNs. To make results comparable, we restricted the analysis to only models trained on ImageNet-1k.
We found that the overall performance is no different (CNNs: $32.49\%$ on average, ViTs: $31.24\%$ on average, t-test of means $p=0.1108$: not significant). Integration biases between transformers and CNNs are comparable: ViT ($7.8\%$) and CNN ($10.1\%$) are not statistically significantly different ($t=2.45,p=0.016$) at the 0.01 criterion.
The way model performance scales across the number of elements is the same across models and across conditions: only one condition across all conditions is different between CNNs and ViTs (the 16-phosphenes condition), while the differences between the rest are statistically non-significant.
Figure 15: Vit vs CNN analysis. (Left): Across all models trained on ImageNet, ViTs and CNN do not exhibit differences across any measures. A: Integration bias. B: Average performance (both conditions). C: Average performance (phosphenes). D: Average performance (segments). (Right): Performance scaling across percentage of elements.